# Robust and Fine-Grained Detection of AI Generated Texts

### v2

@@ Abstract
An ideal detection system for machine generated content is supposed to work well on any generator as many more advanced LLMs come into existence day by day. Existing systems often struggle with accurately identifying AI-generated content over shorter texts. Further, not all texts might be entirely authored by a human or LLM, hence we focused more over partial cases i.e human-LLM co-authored texts. Our paper introduces a set of models built for the task of token classification which are trained on an extensive collection of human-machine co-authored texts, which performed well over texts of unseen domains, unseen generators, texts by non-native speakers and those with adversarial inputs. We also introduce a new dataset of over 2.4M such texts mostly co-authored by several popular proprietary LLMs over 23 languages. We also present findings of our models’ performance over each texts of each domain and generator. Additional findings include comparison of performance against each adversarial method, length of input texts and characteristics of generated texts compared to the original human authored texts.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly narrowed the gap between machine-generated and human-authored text. As LLMs continue to improve in fluency and coherence, the challenge of reliably detecting AI-generated content could become increasingly critical. This issue is particularly pressing in domains such as education and online media, where the authenticity of textual material is paramount. While early efforts such as the GLTR (Gehrmann et al., 2019a) provided valuable insights by leveraging statistical methods to differentiate between human and machine text, these methods often lag behind the rapid pace of LLM evolution. Likewise, initiatives aimed at mitigating neural fake news (Zellers et al., 2019a) have made significant strides in addressing the societal implications of AI-generated misinformation. However, as LLMs become more sophisticated, existing detection systems must be re-evaluated and enhanced to maintain their effectiveness. Further, Each domain comes with its version of the issue of detecting machine generated texts. For instance, proprietary LLMs with internet access and better knowledge cutoffs are more likely to be used in domains like academia. Similarly, bad actors might use an open source generators for the task of creating misinformation and deception through machine generated online content as such models can be hosted locally to not leave a trail and are more flexible in terms of not denying user requests. Hence, tailoring models and approaches for each specific domain/scenario might be better applicable for practical scenarios. We chose a token-classification approach to train a model for the task of distinguish writing styles within a text if more than one were found. This approach helped us achieve better performance over texts of unseen features (i.e domain, generator, adversarial inputs, non-native speakers’ texts) as our models were trained to distinguish different styles within a text rather than classifying an input text as one of the two classes it was trained on. Further, we explored the findings and results upon testing our models over other benchmarks which consist of texts from unseen domain and generators. We also tested our models over benchmarks which consist of texts with various adversarial inputs and those written by non-native speakers. We feel our findings and datasets can aid in further research into mitigating the harms of AI generated texts.
@@ 2 Related Works
A major portion of current research in detecting machine-generated content focuses on longer-form writing through binary classification. However, AI-generated misinformation is more likely to cause harm than its use in academia, making the distinction between AI and human-generated texts on social media platforms a critical challenge. Existing methods often struggle with accurately identifying AI-generated content over shorter texts. Moreover, binary classification approaches, which categorize texts as either human or AI-generated (Wang et al., 2024a), (Wang et al., 2024b), (Bhattacharjee et al., 2023), (Zellers et al., 2019b), (Macko et al., 2023), (Ghosal et al., 2023), (Dugan et al., 2024) are less practical in settings where texts could be co-authored by both humans and LLMs. In contrast, binary classification may be more effective for shorter texts commonly found on reviews and social media platforms (Macko et al., 2024a), (Ignat et al., 2024), where content typically consists of one or two sentences. Additionally, some detection works rely on detecting watermarks from AI-generated texts, (Chang et al., 2024), (Dathathri et al., 2024), (Sadasivan et al., 2024), (Zhao et al., 2023) but not all generators utilize watermarking limiting the applicability of such approaches. Few other approaches utilize statistical methods (Mitchell et al., 2023), (Kumarage et al., 2023), (Gehrmann et al., 2019b), (Hans et al., 2024), (Bao et al., 2023), but they can be prone to mis-classification against adversarial methods like rephrasing and humanizing. (Abassy et al., 2024) introduced a 4-way classification as entirely human authored, entirely llm authored, human-edited and llm-authored or llm-edited human-authored. An ideal detection system should be capable of identifying AI-generated content from any generator without depending on watermarking, especially since watermarking techniques may not be effective for shorter texts. Further an ideal detector should be robust against adversarial methods. To properly deal with co-authored text cases, a token classification approach to detect boundaries (Dugan et al., 2022), (Macko et al., 2024b) between machine authored and human authored portions might be more appropriate. Further in cases of AI usage in scenarios like academic cases, users are likely to use a proprietary LLM with better knowledge cutoffs than an open source LLM. Similarly, for AI misuse over social platforms, users are more likely to use a open-sourced model due to better flexibility and privacy. Hence, building models and benchmarks with a appropriate set of LLMs might be more applicable for practical scenarios. Many proprietary systems struggle at the task of fine-grained detection, further a large enough dataset to cover all POS-tag bi-grams of the text boundaries is required for such fine-grained detectors to work well (Kadiyala, 2024). Previous works in the similar direction include (Lee et al., 2022), (Zhang et al., 2024), (Dugan et al., 2023), (Macko et al., 2024b), (Liang et al., 2024) which utilize a dataset of limited size and limited number of generators or those less likely to be used, which might not be enough for a detector to work well on unseen domains and generators’ texts. Further, the task of detection of such human-llm co-authored texts is a harder task compared to binary classification of texts based on authorship (Geng and Trotta, 2025) (Huang et al., 2025).
@@ 3 Dataset
Our dataset consists of around 2.45M samples. We used 12 different LLMs out of which 9 are popular proprietary LLMs : GPT-o1 (OpenAI, 2024), GPT-4o (etal., 2024), Gemini-1.5-Pro (DeepMind, 2024), Gemini-1.5-Flash, Claude-3.5-Sonnet (Anthropic, 2023), Claude-3.5-Haiku, Perplexity-Sonar-Large (Perplexity, 2023), Amazon-Nova-Pro (Intelligence, 2024), Amazon-Nova-Lite. We also included 3 open-source LLMs i.e Aya-23 (Aryabumi et al., 2024), Command-R-Plus (Cohere For AI, 2024), Mistral-large-2411 (Mistral AI, 2024) which produced outputs that are relatively difficult to distinguish from human written texts compared to other similar models in other benchmarks [^1] https://raid-bench.xyz/ as well as our own datasets. The samples range from 30 to 25K words in length with an average length of around 600 words. Table 1 provides an overview of comparison with other datasets and benchmarks for fine-grained detection of human-llm co-authored texts. Our dataset utilizes a better choice of generators which are more likely to be used in practical scenarios where as prior datasets are limited to smaller and limited number of models. Our dataset also compromises of over 100 times more samples, aiding other researchers in the field.
@@ 3.1 Dataset Distribution
The language distribution of the dataset and LLMs used can be seen in Figure 1. Each language-LLM pair has roughly 10000 samples. Among each set of the 10000 samples; training, development and test sets constitute 40%, 10%, 50% respectively. Additionally, among each set of 10000 samples, 10% were Completely human written, another 10% completely machine generated, and the other 80% were human-LLM co-authored i.e few portions of the text are machine generated and the rest are human written.
@@ 3.2 Dataset Creation
GPT-4o was used through a Azure OpenAI endpoint [^2] https://azure.microsoft.com/en-us/products/ai-services/openai-service/. Command-r-plus and Aya-23 were used through Cohere’s API platform [^3] https://dashboard.cohere.com/. Rest of the models were used through open router’s [^4] https://openrouter.ai/models API. The Rewritten samples were created by providing the generator LLM with the original text and a random prompt among writing an alternate version, a later update of what happened or a rephrased version of the same text. The samples which returned the exact text or a very similar text were once again regenerated. The partially machine generated texts were created by splitting the text at random locations and the generator was asked to finish the text. The split locations were chosen randomly from between the 30th word up to end of text. This was done to provide the LLM with enough context to better work towards text completion.
@@ 3.3 Original Data Source and Filtering
With a goal of training on one domain and testing on every other, we chose to train on old newspapers (HC-Corpora, ) as it has sufficient number of samples i.e 17.2M for 67 languages of the same domain. We then removed samples which originated after release of gpt-3 to avoid mislabelling of samples in our dataset. Further we sampled texts which were at least 3 sentences or 50 words long. For Chinese and Japanese, we sampled texts which were at least 100 characters long.
@@ 4 Our System
We have experimented with various multilingual transformer models (He et al., 2023), (Conneau, 2019), (Beltagy et al., 2020) with/without additional LSTM (Hochreiter, 1997) or CRF layers (Zheng et al., 2015) through a binary token-classification approach. We found that using additional CRF layer produced better results compared to other setups with the same model. All of the transformer models tested have produced nearly identical results over our test set. However, XLM-longformer gave better results over unseen domains and generators’ texts, and was used in the end given the longer default context length of 16384. The token level predictions by the models were then mapped into word-level predictions. We use the model’s predictions to separate text portions based on perceived authorship. Improving the performance required balancing pre and post boundary POS tag bi-grams to reduce error rates (Kadiyala, 2024).
@@ 5 Evaluation and Results
We evaluate the models at 3 levels of granularity : word level, sentence level and overall. For Chinese and Japanese, we performed evaluation at a character level instead of word-level. Each domain and user might have a different preference towards metrics and evaluation, hence we report 3 metrics at each level of granularity : accuracy, recall and precision. For word level mapping of predictions, in cases where part of a word i.e a few tokens are classified differently than others, we assigned the same label to the word as its first token. While mapping word level predictions to a sentence we used majority voting, and in cases where consensus is not obtained, we assigned the same label as the first word. For evaluation over other benchmarks requiring binary classification of texts as human or machine written, we assign a human written label to the text if at least two thirds of the words get classified as human written. We also report several metrics, some of which can be seen in the below tables, rest can be found in Appendix D.
@@ 5.2 Unseen Domains & Unseen Generators
The models were tested twice over (Wang et al., 2024a): once by training on just 10000 samples of a single generator (Aya-23) and again later by training over our complete training data. The benchmark consists of 11,123 samples of peer reviews and student essays (Koike et al., 2024), the generators used were various versions of llama-2 and chat-gpt (earlier version of gpt-4). the samples would hence be from completely unseen domains and generators to our models. The results of both models can be seen in Table 3.
5.3 Unseen Domains & Unseen Generators & Non-Native Speakers
The models were tested by training on just 10k samples each from Aya-23 for English and Arabic Separately. The benchmark’s samples for Arabic were from (Alfaifi, 2013) and (Zaghouani et al., 2024). The samples for English consist of ETS and IELTS student essays sampled from non-native speakers (Chowdhury et al., 2025). Our models were used for inference directly over these texts and the strings of predicted tokens were then used to for binary classification based on how frequently the perceived authorship changed from human to LLM and vice-versa i.e the number of changes and whether the longest string consists of ones or zeroes. The metrics obtained for each language can be seen in Table 4.
@@ 5.4 Unseen Domains & Partially Seen Generators & Adversarial Inputs
We have also tested over raid-bench (Dugan et al., 2025) which consists of texts from 11 generators and 8 domains. among them roughly 10% would be from a seen domain (news articles) while the rest are unseen by our models. The dataset’s texts were also created using various sampling strategies (greedy, random, etc.). The texts were also modified to have adversarial methods including homo-glyphs, mis-spellings, alternative spellings, article deletion etc. Among the 11 generators used, Gpt-4 is one which is similar to the generator whose outputs our model has been trained on (Gpt-4o). However, both of them have different linguistic and stylistic features, similar to how Gpt-4 is different from Gpt-3. We have tested our model’s performance once again upon being trained on our own full training data. Additionally, we have also performed an error analysis to find out what domains, models, attack strategies and decoding strategies effected the model’s performance and to what extent. This can be seen in Figure 2, Figure 3, Figure 4 and Figure 5. The texts were classified as machine generated if at least one third of the tokens within the model’s context length were classified as machine generated. The F1 score obtained with the initial model trained on a single generator was 0.63 and the F1 score grew to 0.79 upon being trained on our full dataset. Evaluation was done directly without performing any preprocessing of the texts and neither were our models trained on texts with any of the adversarial methods.
@@
Figure 3: F1 scores VS the generator’s texts : instruct/chat variants were harder to detect.
@@
5.5 Comparisons with proprietary systems
While many proprietary systems claim to have excellent results, they often struggle with unseen domains and generators. Further many systems like ZeroGPT [^5] https://github.com/zerogpt-net/zerogpt-api and GPTZero [^6] https://gptzero.stoplight.io/docs/gptzero-api/ do not provide fine-grained predictions through their API. This would require manually evaluating the results through the UI by counting correct and incorrectly classified word counts. Hence, these systems were not used for comparisons for fine-grained scores. Several users have tried this manfully over a subset (Kadiyala, 2024) only to find a large gap in performance. For binary classifications as well, they have a minimum length threshold to even work, which would become incompatible over a large number of samples of the benchmark.

### v3

@@ Abstract
An ideal detection system for machine-generated content is supposed to work well on any generator as many more advanced LLMs come into existence day by day. Existing systems often struggle with accurately identifying AI-generated content over shorter texts. Further, not all texts might be entirely authored by a human or LLM; hence, we focused more on partial cases, i.e., human-LLM co-authored texts. Our paper introduces a set of models built for the task of token classification that are trained on an extensive collection of human-machine co-authored texts, which performed well over texts of unseen domains, unseen generators, texts by non-native speakers, and those with adversarial inputs. We also introduce a new dataset of over 2.4M such texts, mostly co-authored by several popular proprietary LLMs in over 23 languages. We also present findings of our models’ performance over texts of each domain and generator. Additional findings include a comparison of performance against each adversarial method, the length of input texts, and the characteristics of generated texts compared to the original human-authored texts.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly narrowed the gap between machine-generated and human-authored text. As LLMs continue to improve in fluency and coherence, the challenge of reliably detecting AI-generated content could become increasingly critical. This issue is particularly pressing in domains such as education and online media, where the authenticity of textual material is paramount. While early efforts such as the GLTR (Gehrmann et al., 2019a) provided valuable insights by leveraging statistical methods to differentiate between human and machine text, these methods often lag behind the rapid pace of LLM evolution. Likewise, initiatives aimed at mitigating neural fake news (Zellers et al., 2019a) have made significant strides in addressing the societal implications of AI-generated misinformation. However, as LLMs become more sophisticated, existing detection systems must be re-evaluated and enhanced to maintain their effectiveness. Further, each domain comes with its version of the issue of detecting machine-generated texts. For instance, proprietary LLMs with internet access and better knowledge cutoffs are more likely to be used in domains like academia. Similarly, bad actors might use an open-source generator for the task of creating misinformation and deception through machine-generated online content, as such models can be hosted locally to not leave a trail and are more flexible in terms of not denying user requests. Hence, tailoring models and approaches for each specific domain/scenario might be more applicable for practical scenarios. We chose a token-classification approach to train a model for the task of distinguishing writing styles within a text if more than one was found. This approach helped us achieve better performance over texts of unseen features (i.e., domain, generator, adversarial inputs, and non-native speakers’ texts) as our models were trained to distinguish different styles within a text rather than classifying an input text as one of the two classes it was trained on. Further, we explored the findings and results upon testing our models over other benchmarks, which consist of texts from unseen domains and generators. We also tested our models over benchmarks, which consist of texts with various adversarial inputs and those written by non-native speakers.
@@ 2 Related Works
A major portion of current research in detecting machine-generated content focuses on longer-form writing through binary classification. However, AI-generated misinformation is more likely to cause harm than its use in academia, making the distinction between AI and human-generated texts on social media platforms a critical challenge. Existing methods often struggle with accurately identifying AI-generated content over shorter texts. Moreover, binary classification approaches, which categorize texts as either human or AI-generated, (Wang et al., 2024a, b; Bhattacharjee et al., 2023; Zellers et al., 2019b; Macko et al., 2023; Ghosal et al., 2023; Dugan et al., 2024) are less practical in settings where texts could be co-authored by both humans and LLMs. In contrast, binary classification may be more effective for shorter texts commonly found on reviews and social media platforms (Macko et al., 2024a; Ignat et al., 2024), where content typically consists of one or two sentences. Additionally, some detection works rely on detecting watermarks from AI-generated texts. (Chang et al., 2024; Dathathri et al., 2024; Sadasivan et al., 2024; Zhao et al., 2023) but not all generators utilize watermarking, limiting the applicability of such approaches. A few other approaches utilize statistical methods (Mitchell et al., 2023; Kumarage et al., 2023; Gehrmann et al., 2019b; Hans et al., 2024; Bao et al., 2023), but they can be prone to misclassification against adversarial methods like rephrasing and humanizing. (Abassy et al., 2024) introduced a 4-way classification as entirely human-authored, entirely LLM-authored, human-edited and LLM-authored, or LLM-edited human-authored. An ideal detection system should be capable of identifying AI-generated content from any generator without depending on watermarking, especially since watermarking techniques may not be effective for shorter texts. Further, an ideal detector should be robust against adversarial methods. To properly deal with co-authored text cases, a token classification approach to detect boundaries (Dugan et al., 2022), (Macko et al., 2024b) between machine-authored and human-authored portions might be more appropriate. Further, in cases of AI usage in scenarios like academic cases, users are likely to use a proprietary LLM with better knowledge cutoffs than an open-source LLM. Similarly, for AI misuse over social platforms, users are more likely to use an open-source model due to better flexibility and privacy. Hence, building models and benchmarks with an appropriate set of LLMs might be more applicable for practical scenarios. Many proprietary systems struggle at the task of fine-grained detection; further, a large enough dataset to cover all POS-tag bi-grams of the text boundaries is required for such fine-grained detectors to work well (Kadiyala, 2024). Previous works in the similar direction include (Lee et al., 2022; Zhang et al., 2024; Dugan et al., 2023; Macko et al., 2024b; Liang et al., 2024; Chen et al., 2023), which utilize a dataset of limited size and limited number of generators or those less likely to be used, which might not be enough for a detector to work well on unseen domains and generators’ texts. Further, the task of detection of such human-LLM co-authored texts is a harder task compared to binary classification of texts based on authorship (Geng and Trotta, 2025; Huang et al., 2025).
@@ 3 Dataset
Our dataset consists of around 2.45 M samples. We used 12 different LLMs, out of which 9 are popular proprietary LLMs: GPT-o1 (OpenAI, 2024), GPT-4o (etal., 2024), Gemini-1.5-Pro (DeepMind, 2024), Gemini-1.5-Flash, Claude-3.5-Sonnet (Anthropic, 2023), Claude-3.5-Haiku, Perplexity-Sonar-Large (Perplexity, 2023), Amazon-Nova-Pro (Intelligence, 2024), and Amazon-Nova-Lite. We also included 3 open-source LLMs, i.e., Aya-23 (Aryabumi et al., 2024), Command-R-Plus (Cohere For AI, 2024), and Mistral-large-2411 (Mistral AI, 2024), which produced outputs that are relatively difficult to distinguish from human-written texts compared to other similar models in other benchmarks [^1] https://raid-bench.xyz/ as well as our own datasets. The samples range from 30 to 25,000 words in length with an average length of around 600 words. Table 1 provides an overview of comparison with other datasets and benchmarks for fine-grained detection of human-LLM co-authored texts. Our dataset utilizes a better choice of generators, which are more likely to be used in practical scenarios, whereas prior datasets are limited to smaller and a limited number of models. Our dataset also comprises the largest collection of human-LLM co-authored texts, aiding other researchers in the field.
@@ 3.1 Dataset Distribution
The language distribution of the dataset and LLMs used can be seen in Figure 1. Each language-LLM pair has roughly 10000 samples. Among each set of the 10000 samples, training, development, and test sets constitute 40%, 10%, and 50%, respectively. Additionally, among each set of 10000 samples, 10% were completely human written, another 10% completely machine generated, and the other 80% were human-LLM co-authored, i.e., a few portions of the text are machine generated and the rest are human written.
@@ 3.2 Dataset Creation
GPT-4o was used through an Azure OpenAI endpoint [^2] https://azure.microsoft.com/en-us/products/ai-services/openai-service/. Command-r-plus and Aya-23 were used through Cohere’s API platform [^3] https://dashboard.cohere.com/. The rest of the models were used through OpenRouter’s [^4] https://openrouter.ai/models API. The rewritten samples were created by providing the generator LLM with the original text and a random prompt among writing an alternate version, a later update of what happened, or a rephrased version of the same text. The samples that returned the exact text or a very similar text were once again regenerated. The partially machine-generated texts were created by splitting the text at random locations, and the generator was asked to finish the text. The split locations were chosen randomly from between the 30th word up to the end of the text. This was done to provide the LLM with enough context to better work towards text completion.
@@ 3.3 Original Data Source and Filtering
With a goal of training on one domain and testing on every other, we chose to train on old newspapers (HC-Corpora, ) as it has a sufficient number of samples, i.e., 17.2 M for 67 languages of the same domain. We then removed samples that originated after the release of GPT-3 to avoid mislabeling of samples in our dataset. Further, we sampled texts that were at least 3 sentences or 50 words long. For Chinese and Japanese, we sampled texts that were at least 100 characters long.
@@ 4 Our System
We have experimented with various multilingual transformer models (He et al., 2023), (Conneau, 2019), (Beltagy et al., 2020) with/without additional LSTM (Hochreiter, 1997) or CRF layers (Zheng et al., 2015) through a binary token-classification approach. We found that using an additional CRF layer produced better results compared to other setups with the same model. All of the transformer models tested have produced nearly identical results over our test set. However, XLM-Longformer gave better results over unseen domains and generators’ texts and was used in the end given the longer default context length of 16384. The token-level predictions by the models were then mapped into word-level predictions. We use the model’s predictions to separate text portions based on perceived authorship. Improving the performance required balancing pre- and post-boundary POS tag bi-grams to reduce error rates (Kadiyala, 2024).
@@ 5 Evaluation and Results
We evaluate the models at 3 levels of granularity: word level, sentence level, and overall. For Chinese and Japanese, we performed evaluation at a character level instead of a word level.  [^5] Unlike English, Chinese and Japanese are written without spaces between words, so segmenting text into words requires an external lexicon or trained segmenter and can introduce alignment errors. Using characters as the atomic units avoids these segmentation ambiguities and ensures every unit is evaluated consistently. Each domain and user might have a different preference towards metrics and evaluation; hence, we report 3 metrics at each level of granularity: accuracy, recall, and precision. For word-level mapping of predictions, in cases where part of a word, i.e., a few tokens, are classified differently than others, we assign the same label to the word as its first token. While mapping word-level predictions to a sentence, we used majority voting, and in cases where consensus was not obtained, we assigned the same label as the first word. For evaluation over other benchmarks requiring binary classification of texts as human or machine written, we assign a human-written label to the text if at least one percent of the words get classified as human-written. We also report several metrics, some of which can be seen in the below tables; the rest can be found in Appendix D.
@@ 5.2 Unseen Domains & Unseen Generators
The models were tested twice over (Wang et al., 2024a): once by training on just 10000 samples of a single generator (Aya-23) and again later by training over our complete training data. The benchmark consists of 11,123 samples of peer reviews and student essays (Koike et al., 2024); the generators used were various versions of Llama-2 and ChatGPT (an earlier version of GPT-4). The samples would hence be from completely unseen domains and generators to our models. The results of both models can be seen in Table 3.
5.3 Unseen Domains & Unseen Generators & Non-Native Speakers
The models were tested by training on just 10,000 samples each from Aya-23 for English and Arabic separately. The benchmark’s samples for Arabic were from (Alfaifi, 2013) and (Zaghouani et al., 2024). The samples for English consist of ETS and IELTS student essays sampled from non-native speakers (Chowdhury et al., 2025). Our models were used for inference directly over these texts, and the strings of predicted tokens were then used for binary classification based on how frequently the perceived authorship changed from human to LLM and vice versa, i.e., the number of changes and whether the longest string consists of ones or zeroes. The metrics obtained for each language can be seen in Table 4.
System Source M E R Avg
GLTR (Gehrmann et al., 2019b) 0.706 0.943 0.773 0.807
Binoculars (Hans et al., 2024) 0.822 0.995 0.613 0.778
RADAR (Gu et al., 2025) 0.768 0.939 0.716 0.794
Current work Our Baselines 0.713 0.914 0.811 0.812
Table 5: Comparisons with other systems on the same benchmarks: MGTD-Bench (M), Academic-Essays (E), and Raid-Bench (R) through F1 Scores, and the average scores from the 3 benchmarks (AVG). The results cover a randomly sampled subset of each benchmark spanning over 13k samples, including those with adversarial methods (R) and non-native speakers (E). Randomly sampled subsets were used with equal distribution of all generators, domains, and adversarial methods distribution in the subset for each benchmark.
@@ 5.4 Unseen Domains & Partially Seen Generators & Adversarial Inputs
We have also tested over raid-bench (Dugan et al., 2025), which consists of texts from 11 generators and 8 domains. Among them, roughly 10% would be from a seen domain (news articles), while the rest are unseen by our models. The dataset’s texts were also created using various sampling strategies (greedy, random, etc.). The texts were also modified to have adversarial methods, including homoglyphs, misspellings, alternative spellings, article deletion, etc. Among the 11 generators used, GPT-4 is one that is similar to the generator whose outputs our model has been trained on (GPT-4o). However, both of them have different linguistic and stylistic features, similar to how GPT-4 is different from GPT-3. We have tested our model’s performance once again upon being trained on our own full training data. Additionally, we have also performed an error analysis to find out what domains, models, attack strategies, and decoding strategies affected the model’s performance and to what extent. This can be seen in Figure 2, Figure 3, Figure 4, and Figure 5. The texts were classified as machine generated if at least one percent of the tokens within the model’s context length were classified as machine generated. The F1 score obtained with the initial model trained on a single generator was 0.63, and the F1 score grew to 0.79 upon being trained on our full dataset. Evaluation was done directly without performing any preprocessing of the texts, and neither were our models trained on texts with any of the adversarial methods.
@@
Figure 3: F1 scores VS the generator’s texts : instruct/chat variants were harder to detect than the corresponding base models.
@@
5.5 Comparisons with other systems
While many proprietary systems claim to have excellent results, they often struggle with unseen domains and generators. Further, many systems like ZeroGPT [^6] https://github.com/zerogpt-net/zerogpt-api and GPTZero [^7] https://gptzero.stoplight.io/docs/gptzero-api/ do not provide fine-grained predictions through their API. This would require manually evaluating the results through the UI by counting correct and incorrectly classified word counts. Several users have tried this manually over a subset (Kadiyala, 2024) only to find a large gap in performance. However, for providing a comparison, we use other open-source systems for binary classification over the same benchmarks.